1. Calculate the percent composition by mass of nitric acid.

2. What is the percent composition of chromium in K2Cr2O7?

3. Find the following:
Given: Sodium Dihydrogen Phosphate, NaH2PO4

a.) formula weight
b.) percent composition of oxygen
c.) weight in grams of 2.7 moles
1
Expert's answer
2019-11-11T07:48:08-0500

Solution.

"w(E) = \frac{Ar \times n \times 100}{M}"

1.

"w(H) = \frac{1*1*100}{63} = 1.59 \%"

"w(N) = \frac{14*1*100}{63} = 22.22 \%"

"w(O) = 100.00 \% - 22.22 \% - 1.59 \% = 76.19 \%"

2.

"w(Cr) = \frac{2 \times 52 \times 100}{294} = 35.37 \%"

3.

a)

"M(NaH2PO4) = 23 + (2 \times 1) + 31 + (4 \times 16) = 120 \ g/mol"

b)

"w(O) = \frac{4 \times 16 \times 100}{120} = 53.33 \%"

c)

"n = \frac{m}{M}"

"m = n \times M"

"m = 2.7 \times 120 = 324.0 \ g"

Answer:

1.

w(H) = 1.59 %

w(N) = 22.22 %

w(O) = 76.19 %

2.

w(Cr) = 35.37 %

3.

a)

M(NaH2PO4) = 120 g/mol

b)

w(O) = 53.33 %

c)

m (NaH2PO4) = 324 g